OpenStudy (i_love_my_nieces):

help fan +medal The sides of a square are increased by a scale factor of 7. The perimeter of the smaller square is 16 ft. What is the perimeter of the larger square?

OpenStudy (anonymous):

Hi, Do you know what a perimeter is ?

OpenStudy (i_love_my_nieces):

A perimeter is the continuous line forming the boundary of a closed geometric figure.

OpenStudy (anonymous):

Yep. It can be viewed as the sum of the length of each of the sides of a geometric figure.

OpenStudy (anonymous):

In a square, the 4 sides are equal. It means that if one side has length \(x\) ft, then there will be 4 sides with length of \(x\).

OpenStudy (anonymous):



OpenStudy (i_love_my_nieces):

So, 4 ft?

OpenStudy (anonymous):

Nope. It will be: \(\large{x + x +x +x = perimeter}\)

OpenStudy (anonymous):

You're told that the perimeter is 16ft. And so: \(\large{x + x +x +x = 16}\) \(\large{4x = 16}\)

OpenStudy (i_love_my_nieces):

My answer choices are: 4 ft. 28 ft. 56 ft. 112 ft.

OpenStudy (anonymous):

Excellent. If you solve for x, you'll have the length of one side. And you're told that each side was stretched by a factor of 7. 'By a factor of something' means that each sides was multiplied by that something.

OpenStudy (anonymous):

Do you guess what would be the next step then?

OpenStudy (i_love_my_nieces):

No.

OpenStudy (i_love_my_nieces):

Is it 28 ft?

OpenStudy (anonymous):

Let find the length of one side. \(\large{4x = 16}\) \(\large{x = \frac{16}{4}}=4\)

OpenStudy (anonymous):

4 was stretched by a factor of 7, which is: \(\large{4*7 = 28}\)

OpenStudy (i_love_my_nieces):

Was I right?

OpenStudy (anonymous):

Your new square looks like this: 

OpenStudy (i_love_my_nieces):

Okay, so, I was right.

OpenStudy (anonymous):

Yep. BUT 28ft is just one side. And the perimeter is the sum of all sides of the square.

OpenStudy (anonymous):

You need to sum four 28's.

OpenStudy (anonymous):

Or: \(4* 28\)

OpenStudy (i_love_my_nieces):

So the perimeter is 112. So what is the answer 28 or 112???

OpenStudy (anonymous):

Since the perimeter is the sum of all sides. The answer is \(4* 28= 112\)ft.